I do not understand this?
Express 50(ten) in binary, trinary and base 4.
thankyou Danielle
: I do not understand this?
Express 50(ten) in binary, trinary and base 4.
thankyou Danielle

Answer by scott8148(2724)

start by extracting the highest power of the base and then work your way down to zero power

50=32+16+2 __ in binary __ 1(2^5)+1(2^4)+0(2^3)+0(2^2)+1(2^1)+0(2^0) __ 110010

50=27+18+3+2 __ in trinary __ 1(3^3)+2(3^2)+1(3^1)+2(3^0) __ 1212

50=48+2 __ in base4 __ 3(4^2)+0(4^1)+2(4^0) __ 302